for the following pair of equations determine algebraically the equilibrium price and output Q=6-p, QS=3P-2


1
Expert's answer
2022-02-28T08:31:43-0500

Q=6-P, QS=3P-2

6-p=3P-2

6+2=3P+P

8=4P

P=8/4

P=2

Q=6-2

Q=4

examination

Q=6-P

Q=6-2=4

QS=3*2-2=4